OpenStudy (anonymous):

Use the distributive property to simplify the expressions below: 8x-10 _____ -2 -3x-9 _______ -3 1/2(g-8)

OpenStudy (anonymous):

do you know the distributive property?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

this is the distributive property: \(\large a(b+c)=a\cdot b + a \cdot c \)

OpenStudy (anonymous):

so the first problem can actually be written as: \(\large \frac{8x-10}{-2} \Rightarrow -\frac{1}{2}(8x-10) \) can you use the distributive property on this expression now?

OpenStudy (anonymous):

i got it